OpenStudy (anonymous):

Let g(x)=x^3+ax+b, where a!=b. If the tangent lines to g at x=a and x=b are parallel, find g(1).

OpenStudy (anonymous):

oh sorry

OpenStudy (loser66):

is it not that g(1) = 1+a +b?

OpenStudy (anonymous):

sure, but I'm looking for a constant answer.

OpenStudy (loser66):

I don't see the link between the condition with the answer :( g'(x) = 3x^2+a so at x =a , g' (a) = 3a^2 +a and g'(b) = 3b^2 +a the tangent lines parallel shows 3a^2+a= 3b^2+a or a =\(\pm b\) but it doesn't affect g(1) . Ha!!

OpenStudy (anonymous):

If it helps, the answer is 1.

OpenStudy (loser66):

if a = -b, then g(1) = 1 but we still have a =b,

OpenStudy (loser66):

@dumbcow

OpenStudy (dumbcow):

@Loser66 , you are correct and it is given that a does not equal b so only option is that a = -b

OpenStudy (loser66):

oops, hehehe...

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

Nice catch. Thanks, guys.

OpenStudy (loser66):

thank you. I forgot the condition.